OpenStudy (anonymous):

when factorised 5(a+b)-9x(a+b) is

OpenStudy (mayankdevnani):

\[\huge \color{green}{WELCOME \space TO \space OPEN \space STUDY}\] @Toby.E

OpenStudy (mayankdevnani):

\[\huge 5(a+b)=5a+5b\] \[\huge -9x(a+b)=-9xa-9xb\] Then, combine all ,we get \[\huge 5a+5b-9xa-9xb\]

OpenStudy (anonymous):

Thanks is that as far as we can solve this?

OpenStudy (mayankdevnani):

no, here is answer \[\huge 5a-9xa+5b-9xb\] \[\huge a(5-9x)+b(5-9x)\] \[\huge (5-9x)(a+b)=\color{red}{answer}\]

OpenStudy (mayankdevnani):

ok.. @Toby.E

OpenStudy (anonymous):

THANKS

OpenStudy (mayankdevnani):

welcome